Milberg Co. uses absorption costing and standard costing to improve cost control.

In 2016, the total budgeted overhead rate was $1.55 per direct labour hour. When preparing the budget, Milberg expected a monthly activity level of 10,000 direct labour hours. The monthly variable overhead cost budgeted for this level of activity was $9,500.

The following data on actual results are provided for the month of November 2016.Materials purchased 20,000 units Direct labour costs incurred $36,000 Total of direct labour rate and efficiency variances 50

Solve for:

Variable Overhead Spending Variance

Variable Overhead Efficiency Variance

Variable Overhead Total Variance

Fixed Overhead Spending Variance

Fixed Overhead Efficiency Variance

Fixed Overhead Total Variance

* please include all formulas and do not use excel

* I have already solved the other variances. Just struggling with the overhead ones
Answer #1
std variable overhead per hour = $9500/10000 hrs=0.95 per hour
Budgeted Fixed Overhead = (1.55-.95)*10000 hours = $6000
Total Labor Variance = (Std hours * Sta Rate ) - (actual Hr *Actul Rate)
500 (F) = (std hrs * $5 per hr) - ([$36000/$4.8] *$4.8)
500 (F) = (std hrs* 5 per hr) -(7500 hrs*4.8 per hr)
500 (F) = (std Hrs *5 per hr)-$36000
Std hrs =7300 hrs
1) Variable Overhead Efficiency Varianace = Std Rate * (std hrs - actual hrs)
= 0.95 per hr * (7300 hrs -7500 hrs)
= 190 (F)
2) Under Applied Variable Overhead Cost = Variable Overhead Spending Variance + Variable Overhead Efficiency Variance
1065 (U) = Spending Variance + (-190) (F)
Therefore variable Overhead Spending Variance = 1065+190 = 1255 (U)
3) Fixed Overhead Efficiency Variance = std rate per hr * ( std hrs - actual hrs)
= 0.6 per hr * (7300-7500 hrs)
=120 (F)
4) total fixed overhead cost variance = total under applied overheads - total variable overhead cost variance
= $2256 (U) - 1065 (U)
=1191 (U)
5) Fixed Overhead spending variance = Total fixed cost variance - fixed overhead efficiency variance
=1191(U) - (-120)(F)
=1311 (U)